Subject 1has the following values: Minute ventilation = 7 liters/minute, Respiratory frequency = 16 Br/minute, PaCO2= 35 mm Hg, PECO2= 27 mm Hg. Identify whether each of the following are correct or incorrect.

a. VT= 2.3 liters

b. VD/VT= 0.23

c.the alveolar volume is 77% of the Vt

d.VA= 5.4 liters/minute
Answer #1

Here Minute Ventilation is 7 liters(7000ml)

So, VT(Tidal Volume Will be) - 7000ml/16 breath/min = 438ml

Option A is incorrect

VD/VT ratio = (PaCO2 - PECO2)/ PaCO2 = (35 - 27) / 35 = 0.228 =0.23

Option B is Correct

Alveolar Volume is = (VT - VD) * RR

VD(Dead space volume) = VT * (PaCO2 - PECO2)/ PaCO2

438ml *0.23 = 101ml So, VD is 101 ml

Alveolar Volume will be (VT - VD) * RR , (438ml - 101ml) *16 = 5392ml/min = 5.4 lt/min

5.4lt its not 77% of vt(438 ml), 5.4lt is 77% of minute volume i.e 7lt

Option C is incorrect

VA(Alveolar volume) is 5.4lt/min as we just calculated.

Option D is correct